OpenStudy (anonymous):

simplify each expression. use positive exponents. (x–2y–4x3) –2

OpenStudy (anonymous):

is the 4x3 4 times 3

OpenStudy (csubero):

yea it is

OpenStudy (anonymous):

-2x+4y+24

OpenStudy (anonymous):

wait so 48*3?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

you cant combine unlike terms

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

wait im coming

OpenStudy (anonymous):

(x−2y−4x3)−2 Whenever you have a negative exponent, you can write it as the reciprocal of the base. This is usually stated this way: a−n=1an We can use this to write (x−2y−4x3)−2=1(x−2y−4x3)2 hope i helped

OpenStudy (anonymous):

y8x2 is ur answer

OpenStudy (anonymous):

(x1y−4)−2=x−2y8 No negative exponents, so we use the property from my first response at the very beginning. x−2=1x2 This gives us x−2y8=y8x2 So your answer is y8x2

OpenStudy (anonymous):

Let's go back to your problem for a moment. (x−2y−4x3)−2 Ignore the fact that there are negative exponents, and ignore the fact that there is a "y" in there. See those two "x" parts? Try combining those together. Simplify x−2x3

OpenStudy (anonymous):

x-5 Ah, when you add -2 and 3, you get 1. That will give you x^1

OpenStudy (anonymous):

are u here

OpenStudy (anonymous):

Okay. Now we have (x1y−4)−2 Let me say a couple things about this. First, these problems are going to be easier if you write the exponents that usually aren't written in there. For example, rewrite x as "x^1" like I did. Second, you can't combine an x and a y. For example, what I'm about to do is WRONG: x1y−4=xy−3 It just simply doesn't work like that. You can ONLY combine them if they have the same BASE.

OpenStudy (anonymous):

Now, let's talk about raising a term to a power. Let's say we have (a^3)^4: (a3)4=a3a3a3a3 Now each of these has the same base, so we can ADD the exponents. (a3)4=a3+3+3+3=a12 New Rule: If we raise a power to a power, we MULTIPLY the exponents. Examples: (a5)4=a20 (g9)10=g90 (z−2)−2=z4 Also, notice the difference between these two: a4a5=a9 (a4)5=a20

OpenStudy (anonymous):

so ur anser is y8x2

OpenStudy (anonymous):

ur welcome